The company is generating a cash flow of $300,000 every year. The market interest rate is 18%. The company has 100,000 shares outstanding. What is the value of the company’s share?
Answer #1

Here cash flow of 300000 is generated for infinite period

Now cash flow per share is 3(300000/100000)

So present value formulae is derived as below

Pv =3/1.18 + 3/(1.18)^2---------------3/(1.18)^infinity

= 3( sum of infinite terms of geometric progression) = 3(a)/r-1

= 3( 1/1.18)/0.18/1.18

= 3/0.18 = 16.67

So share value is 16.67

The above process is derivation of formula

You can directly use formulae which is cash flow / rateof interest